OpenStudy (anonymous):

I understand to solve: If the perimeter of a rectangular property is 46 meters, and the area of the property is 76 meters, what is the length of each of the shorter sides you will get to 23y-y^2-76=0 BUT how do you get from there to the answer Length is 19, Width is 4?

OpenStudy (anonymous):

Rewrite the equation to be y^2-23y+76=0, then factor to (y-19)(y-4)=0 and solve by setting each binomial to zero and you get the 19 and 4. Note on factoring: (-19)(-4)=76 and 19+(-4)=-23.